OpenStudy (anonymous):

choose an equation for the line that is parallel to the given line for passes through the given point. y = 5x + 10; (2, 14) a. y = 1/5x + 4 b. y = -1/5x - 4 c. y = 5x - 68 d. y = 5x + 4

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

y = mx+c m = slope of the line y = 5x + 10 m = 5 two parallel lines have same slope using point slope formula y - 14 = 5(x - 2) make this in std form of y = mx+c